calculus 2
5. Determine the volume of the solid formed by revolving the region bounded by y = cos(x), y = 0, x = 0 about x = 7/2
Answer #1


5 g = cos(a), y=0, 2=0, x=R Shell method b . vowme V= J 217(V2-2) (9)d1 19. my=07 0 = Color x= 22 VE S 20 ( 02-a( cofx) dy V=